OpenStudy (andiemarinn):

How do you start this??? What is the closed form of a geometric sequence with ratio between consecutive terms 5 and inital term 10? A. an = 5(2n) B. an = 2(5n) C. an = 5n + 5 D. an = 5n

OpenStudy (phi):

ratio between consecutive terms 5 means each term is 5 times bigger than the one before it. inital term 10 means it starts at 10 so the sequence is 10, 50, 250, 1250, ...

OpenStudy (phi):

I would try each formula with n=1 and n=2 and see if I get 10 for n=1, and 50 for n=2

OpenStudy (andiemarinn):

okay thank you so much!

OpenStudy (phi):

the formula http://www.mathsisfun.com/algebra/sequences-sums-geometric.html you would use \[ a r^{n-1} \] with a=10, r= 5 \[ 10 5^{n-1} \] we can change its form by doing this: \[10\cdot 5^{n-1}= 10\cdot 5^n \cdot 5^{-1} \\ = 10 \cdot 5^n \cdot \frac{1}{5^1} \\= \frac{10}{5} \cdot 5^n \\= 2 \cdot 5^n \]

OpenStudy (jtug6):

ohhhh. just simplifying with algebra. i shouldve tried that. LOL.